1. In a hydroelectric power plant, water enters the turbine nozzles at 800 kPa absolute with a low velocity. If the nozzle outlets am exposed to atmospheric pressure of 100 kPa. determine the maximum velocity to which water can be accelerated by the nozzles before striking the turbine blades.



2. A Pitot-static probe is used to measure the speed of an aircraft flying at 3000 m. If the differential pressure reading is 3 kPa, determine the speed of the aircraft



3. A piezometer and a Pitot tube are tapped into a 4-cm-diameter horizontal water pipe, and the height of the water columns are measured to be 26 cm in the piezometer and 35 cm in the Pitot tube (both measured from the top surface of the pipe). Determine the velocity at the center of the pipe.


1
Expert's answer
2021-12-15T16:22:24-0500

1.

"v=\sqrt{\frac{2\Delta p}{\rho}}=40~\frac ms."

2.

"v=\sqrt{\frac{pg}{2h}}=108~\frac{km}h."

3.

"v=\sqrt{\frac{dh}{\rho g(h+s)}}=0.1~\frac ms."